OpenStudy (anonymous):

11.) Refer to right triangle. If m Ang A = 30° and BC = 5, what is the length of AC? 5 sqrt(2) 5 5 sqrt(3) 10

OpenStudy (anonymous):

Angle B is righ angl?

OpenStudy (anonymous):

I do B is righ angle than sin30=5/AC AC=5/sin30 Ac=10